# WarriorCoder: Learning from Expert Battles to Augment Code Large Language Models

### v1

@@ 4.3 Ablation Study
From HumanEval HumanEval+ MBPP MBPP+
Qwen 67.7 60.4 68.3 55.8
Mistral 70.1 64.0 67.7 54.8
All (ours) 72.1 66.0 70.5 57.8
Table 2: The results observed when learning from different teachers.
Table 2 presents the results observed when the target model learning from different teachers. Qwen and Mistral perform well on the HumanEval (HumanEval+) and MBPP (MBPP+) benchmarks, respectively. This indicates that each of these models enables the target model to learn a specific set of knowledge. However, WarriorCoder benefits from learning across all expert code LLMs, leveraging the strengths of each. As a result, it outperforms the others across all four benchmarks, demonstrating the advantages of integrating knowledge from multiple specialized teachers.
@@ 4.4.1 Diversity Analysis
Figure 3 illustrates the overlap between the instructions mined from expert LLMs and those from existing training datasets, measured using the ROUGE score. The majority of the mined instructions exhibit a ROUGE score of less than 0.3, indicating that they are largely distinct from the instructions in these datasets. Notably, there are no mined instructions with a ROUGE score exceeding 0.6, further reinforcing the observation that the mined instructions are sampled from the internal distribution of expert LLMs but not simply replications or extensions of existing training data. As a result, these instructions demonstrate a higher degree of diversity, making them particularly valuable for training purposes, as they introduce novel examples that can enhance the capability of the target model.
@@ 4.4.2 Difficulty Analysis
Figure 4 illustrates the difficulty distribution of the mined instructions, providing insights into the experts’ internal knowledge. The majority of the instructions fall within the ’good’ level, with scores ranging from 6 to 8. Instructions rated at the ’excellent’ level (scores 9-10) represent only a small fraction of the total, suggesting that highly complex or advanced tasks are relatively rare. On the other hand, instructions with scores below 6 are excluded from the training set, as they tend to be either too easy or overly ambiguous. Such instructions are considered detrimental to the training stage, as they may not provide meaningful learning signals and could undermine the model’s performance and generalization ability.
Figure 4: The difficulty of mined instructions. As mentioned in Section 3.1, the difficulty of instructions is divided into four levels: excellent (9-10), good(6-8), average(3-5) and poor(1-2).
4.4.3 Case Study
@@
Both instructions are clear and specific, providing a solid foundation for tackling the problem. However, the first instruction is more straightforward, with fewer requirements and a greater emphasis on practical tips and strategies to handle the task. In contrast, the second instruction introduces more complex conditions and does not include any guidance, making it more challenging to solve the problem.
5 Conclusion
This paper highlights the significant limitations of existing data flywheels for code models, which primarily rely on pre-existing datasets and annotations from a limited pool of proprietary LLMs. This dependency leads to a lack of data diversity and reinforces thr systemic biases. Even more concerning is the fact that many current open-source expert code LLMs keep their training data proprietary, further restricting access to diverse and high-quality data sources. To address these challenges, we propose WarriorCoder, which learns from the competitive knowledge of expert code LLMs, enabling the absorption of each expert’s strengths. Unlike existing methods that expand and refine datasets, our approach constructs training data from scratch, leveraging the expertise of the strongest models in a competitive setup. This allows us to achieve competitive performance without the need for pre-existing datasets and costly annotations. Furthermore, our approach is not limited to enhancing LLMs’ coding abilities and can potentially be applied to other complex tasks in the future.

### v3

@@ 4.3 Ablation Study
#Num HumanEval HumanEval+ MBPP MBPP+
1 75.4 72.6 73.3 62.4
2 77.2 73.3 74.5 62.9
5 80.5 75.6 76.2 64.8
Table 5: The results observed when learning from varying numbers of experts.
Table 5 presents the results observed when the target model learns from varying numbers of experts. The target model shows a significant improvement when learning from just one code LLM, indicating that even a single code expert enables it to acquire a specific set of knowledge. However, as the number of experts increases, WarriorCoder benefits from learning across all expert code LLMs. As a result, the model trained with 5 code LLMs outperforms others across all four benchmarks, demonstrating the advantages of integrating knowledge from multiple specialized experts.
@@ 4.4 Data Analysis
4.4.1 Dependence Analysis
Figure 3 illustrates the overlap between the instructions mined from expert LLMs and those from widely used code training datasets, measured using the ROUGE score. The majority of the mined instructions have a ROUGE score of less than 0.3, suggesting they are largely distinct from those in existing datasets. Notably, no mined instructions exceed a ROUGE score of 0.6, further emphasizing that the mined instructions are drawn from the internal distribution of expert LLMs, rather than being simple replications or extensions of the training data. Consequently, these instructions exhibit a higher degree of independence, making them particularly valuable for training, as they provide novel examples that can enhance the target model’s capabilities.
@@ 4.4.2 Diversity Analysis
Figure 4: The heatmap of win rates of the selected code experts.
Table 4 reveals the distribution of different tasks in the training data. The range of instructions covers a variety of tasks, ensuring that WarriorCoder can generalize effectively across multiple benchmarks. Notably, while Code Reasoning represents only 2.9% of the entire dataset, WarriorCoder still achieves outstanding performance on CRUXEval, highlighting the potential of the framework that learns from expert battles. Furthermore, Figure 4 illustrates the battle results between the five selected code experts. Even though an expert may have the highest Elo Rating, it is not necessarily the best performer on all tasks. However, WarriorCoder learns from the winner of each instruction, thereby diversifying the target responses.
@@ 4.4.3 Difficulty Analysis
Figure 5: The proportion of difficulties of mined instructions. As mentioned in Section 3.2, the difficulties of instructions are divided into four levels: excellent (9-10), good(6-8), average(3-5) and poor(1-2).
Figure 5 shows the difficulty distribution of the mined instructions, offering insights into the internal knowledge of the code experts. Most instructions fall within the ’good’ level, with scores between 6 and 8. Instructions rated as ’excellent’ (scores 9-10) constitute only a small portion of the dataset, indicating that highly complex or advanced tasks are relatively rare. Instructions with scores below 6 are excluded from the training set, as they tend to be either too easy or overly ambiguous. Such instructions are considered detrimental to the training stage, as they may not provide meaningful learning signals and could undermine the model’s performance and generalization ability. More examples are listed in Appendix A.
5 Conclusion
This paper highlights the limitations of existing data flywheels for code LLMs that primarily rely on pre-existing datasets and annotations from a limited pool of proprietary LLMs, leading to a lack of data diversity and reinforces the systemic biases. Even more concerning is the fact that many current open-source expert code LLMs keep their training data proprietary, further restricting access to diverse and high-quality data sources. To address these challenges, we propose WarriorCoder, which learns from expert battles, enabling the absorption of each expert’s strengths. Unlike existing methods that expand and refine datasets, we construct training data from scratch and achieve SOTA performances on multiple benchmarks without the need for pre-existing datasets and costly annotations. Furthermore, our approach can potentially be applied to other complex tasks besides coding in the future.
Limitations
In this paper, we propose a novel training paradigm in which the target model learns from expert battles, aiming to overcome the limitations of current data flywheels. While we can generate high-quality and diverse data from scratch at a low cost, the battle process can become time-consuming when the number of experts is large. Exploring more efficient and effective competition modes is a promising direction for future work.
Appendix A Case Study
@@
You are an AI assistant designed to provide helpful, step-by-step guidance on coding problems. The user will ask you a wide range of Python coding questions. Your purpose is to assist users in understanding coding concepts, working through code, and arriving at the correct solutions.
Table 6: The prompt for instruction mining.
This is a chatbot arena. You will be given assistant A’s answer, and assistant B’s answer. Please act as an impartial judge and evaluate the capability of two AI assistants. You should choose the assistant that follows instructions and answers questions better. Your evaluation should consider factors such as helpfulness, relevance, and accuracy. Begin your evaluation by comparing the responses of the two assistants and provide a short explanation. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. DO NOT allow the LENGTH of the responses to influence your evaluation, choose the one that is straight-to-the-point instead of unnecessarily verbose. When the two candidates perform equally well, choose the SHORTER answer. Do not favor certain names of the assistants. Be as objective as possible. After providing your explanation concisely within 200 words, output your final verdict by strictly following this format: "[[A]]" if assistant A is better, "[[B]]" if assistant B is better, and "[[Tie]]" for a tie. Finish your judgement within 300 words.
[[User Question]]
{instruction}
[[The Start of Assistant A’s Answer]]
{#A’s Answer}
[[The End of Assistant A’s Answer]]
[[The Start of Assistant B’s Answer]]
{#B’s Answer}
[[The End of Assistant B’s Answer]]
Table 7: The prompt for the evaluation of pairwise competitions.
Both instructions are clear and specific, providing a solid foundation for tackling the problem. However, the first instruction is more straightforward, with fewer requirements and a greater emphasis on practical tips and strategies to handle the task. In contrast, the second instruction introduces more complex conditions and does not include any guidance, making the problem more challenging to solve.
Appendix B Prompts for Instruction Mining and Evaluation
Table 6 and Table 7 show the prompts for instruction mining and evaluation. To maintain impartiality in the evaluation process, we withhold the names of the opponents from the judges, thus avoiding potential system biases. Additional rules will be discussed in the next section.
Appendix C Rules for the Fairness of Evaluation
To ensure impartiality in the evaluation process, we establish a set of rules, including order shuffling, suspicion averting, and offensive-defense balance.